# Under-coverage in high-statistics counting experiments with finite MC samples

### v2

@@ 3.7 Adding external constraints
In section 3.1, all parameters of the likelihood were assumed to be unconstrained, though this is certainly not the most general case of interest. Indeed, it often happens that at least some of the nuisance parameters can be constrained by independent auxiliary measurements. Let us then assume that the likelihood function includes both constrained and unconstrained nuisance parameters. If the set of $q$ constrained nuisance parameters is denoted by $\boldsymbol{\tau}$, the right-hand side of eq. (25) gets modified to
@@
where the estimators from the auxiliary measurements have been assumed as normally-distributed and independent from each other Pinto:2023yob . The additional parameters $\boldsymbol{\tau}$ can be profiled at a fixed value of $(\mu,\boldsymbol{\theta})$ yielding the well-known result Pinto:2023yob ; Behnke :
$-2\ln\frac{L(\mu,\boldsymbol{\theta},\hat{\boldsymbol{\tau}}_{(\mu,\boldsymbol{\theta}}))}{{L}(\hat{\mu},\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\tau}})}=\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta},\boldsymbol{\tau}_{0})\right)^{T}\left({\bf V}+{\bf K}{\bf K}^{T}\right)^{-1}\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta},\boldsymbol{\tau}_{0})\right),$ (64)
@@
with ${\bf H}={\bf V}^{-\frac{1}{2}}{\bf K}$. The newly introduced matrix ${\bf Z}$ is a symmetric and positive-definite matrix, so it is invertible and has a unique square-root matrix ${\bf Z}^{\frac{1}{2}}$. By means of eq. (3.7), eq. (64) can be ultimately recast in the same form of eq. (26), namely
@@
where ${\bf r}_{\mu}^{\prime}$ and ${\bf A}^{\prime}$ are defined as in eq. (27) with the replacement ${\bf V}^{-1/2}\to\left({\bf V}{\bf Z}\right)^{-1/2}$. In this framework, statistical fluctuations affecting the entries of the Jacobian matrix ${\bf K}$ are propagated into the NLL function by the matrix ${\bf Z}$ which mixes them with fluctuations of ${\bf J}$ in a possibly intricate way. However, due to the formal identity between eq. (26) and (66), the same analysis pursued in section 3.2 still holds. Notice, however, that in the case of eq. (3.7) some of the scaling laws of eq. (55) are broken due to a non-trivial combination of statistical and systematic uncertainties. Also, the inclusion of external constraints in the likelihood function is likely to reduce the global correlation factor of the POI and hence improve the validity of the asymptotic approximation.

### v3

@@ Appendix C Adding external constraints to the likelihood
In section 3.1, all parameters of the likelihood were assumed to be unconstrained, though this is certainly not the most general case of interest. Indeed, it often happens that at least some of the nuisance parameters can be constrained by independent auxiliary measurements.
Let us then assume that the likelihood function includes both constrained and unconstrained nuisance parameters. If the set of $q$ constrained nuisance parameters is denoted by $\boldsymbol{\tau}$, the right-hand side of eq. (28) gets modified to
@@
where the estimators from the auxiliary measurements have been assumed as normally-distributed and independent from each other [23]. The additional parameters $\boldsymbol{\tau}$ can be profiled at a fixed value of $(\mu,\boldsymbol{\theta})$ yielding the well-known result [23, 4]:
$2\ln\frac{L(\mu,\boldsymbol{\theta},\hat{\boldsymbol{\tau}}_{(\mu,\boldsymbol{\theta}}))}{{L}(\hat{\mu},\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\tau}})}=\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta},\boldsymbol{\tau}_{0})\right)^{T}\left({\bf V}+{\bf K}{\bf K}^{T}\right)^{-1}\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta},\boldsymbol{\tau}_{0})\right),$ (60)
@@
with ${\bf H}={\bf V}^{-\frac{1}{2}}{\bf K}$. The newly introduced matrix $\bf Z$} is a symmetric and positive-definite matrix, so it is invertible and has a unique square-root matrix ${\bf Z}^{\frac{1}{2}}$. By means of eq. (C), eq. (60) can be ultimately recast in the same form of eq. (29), namely
@@
where ${\bf r}_{\mu}^{\prime}$ and $\bf A$$^{\prime}$ are defined as in eq. (30) with the replacement ${\bf V}^{-1/2}\to\left({\bf V}{\bf Z}\right)^{-1/2}$. In this framework, statistical fluctuations affecting the entries of the Jacobian matrix $\bf K$} are propagated into the NLL function by the matrix $\bf Z$} which mixes them with fluctuations of $\bf J$} in a possibly intricate way. However, due to the formal identity between eq. (29) and (62), the same analysis pursued in section 3.2 still holds. Notice, however, that in the case of eq. (C) some of the scaling laws of eq. (46) are broken due to a non-trivial combination of statistical and systematic uncertainties. Also, the inclusion of external constraints in the likelihood function is likely to reduce $\rho_{\mu}$ and hence improve the validity of the asymptotic approximation.